# PLM-Based Discrete Diffusion Language Models with Entropy-Adaptive Gibbs Sampling

### v1

@@ D.4 Deontology
D.5 RocStories
@@
Appendix E Details on Text Augmentation Using GPT models
E.1 GPT-3.5turbo ~ GPT-4-Omni
@@ Appendix E Details on Text Augmentation Using GPT models
Appendix F Details on CDLMs
F.1 Experimental Details
@@ F.2 Results Interpretations
Appendix G Details on Translation Results
G.1 Comparison Between Easy-First and Our Proposed Method
@@ G.2 Experimental Details
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix F.1.
SEDD The SEDDLou et al. (2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate dataset-guided generation. To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols. Specifically, the input and output token lengths were constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics. Moreover, distinct special tokens were introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training. Individual data entries were further demarcated by an EOS token to delineate discrete sequences within the training process.
@@
G.3 Experimental Results
G.3.1 NAR Transformer, DisCo
@@ G.3.2 Diffuseq-v2
Appendix H Entropy Flow
@@
In Figure 4, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes. The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process. This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 2, 3, and 4.

### v2

@@ F.2 Deontology
F.3 QQP
Table 18: QQP generation examples
QQP Generation Output Examples of Diffusion-EAGS
Output $X_{2}$ “What are the best known fiction and books of Isaac Asimov?”
Constraint $Y$ “Can we ever store energy produced in lightning?”
Output $X_{1}$ “How do we store heat energy from lightning?”
Output $X_{2}$ “How can you store energy from lightning?”
F.4 QG
Table 19: QG generation examples
QG Generation Output Examples of Diffusion-EAGS
Output $X_{1}$ “What kind of bird can fly backwards?”
Output $X_{2}$ “Which bird is able to fly backwards?”
Output $X_{1}$ “What is a pouch?”
Output $X_{2}$ “What is the smallest animal without a pouch.”
F.5 RocStories
@@
Appendix G Details on Text Augmentation Using GPT models
G.1 GPT-3.5turbo ~ GPT-4-Omni
@@ Appendix G Details on Text Augmentation Using GPT models
Appendix H Details on CDLMs
H.1 Experimental Details
@@ H.2 Results Interpretations
Appendix I Details on Translation Results
I.1 Comparison Between Easy-First and Our Proposed Method
@@ I.2 Experimental Details
Diffuseq-v2 For Diffuseq-v2, we employ the vocabs of mBERT and choose 128 as max length for ende translation. Other settings are identical as in Appendix H.1.
SEDD The SEDDLou et al. (2024) model, originally designed for open-ended text generation, was adapted in this study to facilitate conditional generation. To align the model’s architecture with the specific requirements of the structured dataset, several modifications were implemented in both hyperparameters and preprocessing protocols. Specifically, the input and output token lengths were constrained to a range of 64 to 128 tokens, ensuring a more appropriate fit to the dataset’s structural characteristics. Moreover, distinct special tokens were introduced to clearly differentiate between input and output sequences, thereby enhancing the model’s ability to distinguish between these components during training. Individual data entries were further demarcated by an EOS token to delineate discrete sequences within the training process.
@@
I.3 Experimental Results
I.3.1 NAR Transformer, DisCo
@@ I.3.2 Diffuseq-v2
Appendix J Entropy Flow
@@
In Figure 4, we illustrate the tendency of the sequential sum of entropy for various discrete generation processes. The changes of entropy during the generation process in Diffusion-EAGS, represented by the yellow line, show that our model effectively follows a gradual decrease in entropy, mirroring the inverse trend of the training process. This gradual change in entropy facilitates successful DDLM training, which results in superior text quality performance compared to other diffusion models, as demonstrated in Tables 3, 2, and 4.